Find the section of the conicoid x^2/2-y^2/3 = 2z by the plane x-2y+z = 1. What conic does this section represent? Justify your answer.

1
Expert's answer
2021-11-01T16:20:24-0400
"x-2y+z=1=>z=-x+2y+1"

"\dfrac{x^2}{2}-\dfrac{y^2}{3} = 2(-x+2y+1)"

"\dfrac{x^2+4x+4}{2}-\dfrac{y^2+12y+36}{3} =2+2-12"

"\dfrac{(y+6)^2}{24}-\dfrac{(x+2)^2}{16} =1"

The equation represents the hyperbola.